Determine the exact rectangular coordinates of the point whose coordinates are P[-4,-3pie/4]
Answer by lwsshak3(11628)

Determine the exact rectangular coordinates of the point whose coordinates are P[-4,-3pie/4]
..
(-4,-3π/4)
r=-4
angle a=-3π/4
reference angle=π/4 in quadrant III where sin and cos are negative
x=r cos a=-4(-√2/2)=2√2
y=r sin a=-4(-√2/2)=2√2
ans:
rectangular coordinates at P: (2√2,2√2)